OpenStudy (anonymous):

jake used an employment agency to find him a new job. jake has agreed to pay the agency a commission equal to 4% of his first years salary. if the agency finds jake a job that pays $7,500 a month , what amount of commission will jake pay?

OpenStudy (anonymous):

How many months are there in a year?

OpenStudy (anonymous):

@mcnabkt

OpenStudy (anonymous):

well it says something about 4%

OpenStudy (anonymous):

I said "How many months are there in a year?"

OpenStudy (mathmath333):

if jake's salary for one month is \($ ~7500\) what will be his salary for one year

OpenStudy (anonymous):

okay so i divide that by 4% right?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Ugh....

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

You multiply it to how many months there are in a year...

OpenStudy (anonymous):

So how many months are there in a year?

OpenStudy (anonymous):

is it 7500 * 12/4% ?

OpenStudy (anonymous):

No. Just do what I tell you to do.

OpenStudy (anonymous):

So there are 12 months in a year, so multiply that to 7500. 7500 x 12 = ?

OpenStudy (anonymous):

@mcnabkt

OpenStudy (anonymous):

90000

OpenStudy (anonymous):

now divide by 4%?

OpenStudy (anonymous):

No, you multiply 4% to 90000

OpenStudy (anonymous):

But we need to change it to a decimal or a fraction.

OpenStudy (anonymous):

4% is the same as 0.04, so what is 0.04 x 90000? Or, you could do 4/100 x 90000.

OpenStudy (mathmath333):

\(90000\times 4\%\) u have to multily it

OpenStudy (anonymous):

\(0.04\ \times\ 90000\) \(\dfrac{4}{100}\times\ 90000\)

OpenStudy (anonymous):

You can do one of those, they'll give you the same answer.

OpenStudy (anonymous):

so 3600?

OpenStudy (anonymous):

Correct! :)

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

?

OpenStudy (anonymous):

because i have more questions hahahaha

OpenStudy (anonymous):

Ok. :)